if two people comple a job in 40 minutes and it takes one to do it in 70 mins. how long does it take the other?
Answer by ewatrrr(24785)

 
Hi,
if two people comple a job in 40 minutes and it takes one to do it in 70 mins.
how long does it take the other?
Per minute the Equalizer
1/70 + 1/x = 1/40 |Multiplying thru by 40*70*x so as all denominators = 1
40x + 40*70 = 70x
2800 = 30x
x = 93 1/3 min